Grade 12th passMechanics

A man suspended from a spring execute shm when the body is at helf the amplitude far from the mean position the ratio of the kinetic energy to the potential energy of the body will be

Tanu Dey
9 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago
Potential energy at A/2 is:
U = Kx^2/2 => k(A/2)^2/2 = > KA^2/8 J
Kinetic energy at A/2 is:
K = 0.5K(A^2 – x^2) => 0.5k(A^2 – A^2/4) => 3KA^2/8 J
Ratio => K/U = > 3/1
or
3:1